# Escaping the Subprime Trap in Algorithmic Lending

### v1

@@ 2.2.4 Learning through lending
If bank $j$ lends to group $i$, it observes a return $\pi_{i}^{(t)}$ and then updates its posterior beliefs about the variance $\sigma_{i}^{2}$. After $M$ such observations, its posterior belief takes the form of:
$\hat{\sigma}_{i}^{2(M)}=\dfrac{1}{M+1}\left(\hat{\sigma}_{i}^{2(0)}+\sum_{m=1}^{M}\Bigl{(}\pi_{i,B}^{(m)}-\overline{\pi}_{i,B}^{(M)}\Bigr{)}^{2}\right)\quad\textrm{Where: }\quad\overline{\pi}_{i,B}^{(M)}=\frac{1}{M}\sum_{m=1}^{M}\pi_{i,B}^{(m)}$
@@
2.3 Our Modeling Choices
Payoff Distribution We assume returns are normally distributed about some known location $\mu$. Assuming that the means are fixed and known is WLOG, since we only suppose this for the sake of illustrating the role of prior beliefs. It is straightforward to adapt these results to cases with unknown or different group means. Second, our use of the normal distribution reflects the fact that the size of the principal applied for by a given loanee is a random quantity, in addition to their repayment risk. Heuristically, we can decompose $\pi$ into $\pi=\mathbb{P}(\text{Repayment})\times P-\mathbb{P}(\neg\text{Repayment})P$, where $P$, the value of the loan principal, is random. This quantity is then approximately normally distributed. Replacing this with any heavier-tailed distribution would yield analogous conclusions (with different formulas for the variance thresholds).
[…]
@@ 2.3 Our Modeling Choices
Inaccurate prior beliefs Our model supposes that banks have inaccurate beliefs about the creditworthiness of applicants. This is a plausible outcome when decisions are made on the basis of credit scoring data that is either biased or noisy. Credit scores may be inherently noisier representations of underlying default risk for minority groups Blattner and Nelson [2021]. Credit scoring data for minority groups is more likely to be missing Kenneth P. Brevoort [2015], while the components of the score may themselves be low quality indicators of repayment ability Rice and Swesnik [2013].
3 Single-Period Setting
We first study the stage game, to show how the VaR constraint affects lending to group $B$ in each period.
First we assume that Banks have prior beliefs about group variances that satisfy the following conditions.
Assumption 3 (Prior variances).
$\sigma_{W}\leq\dfrac{\rho_{L}-\mu}{\Phi^{-1}(\alpha)}<\hat{\sigma}^{(0)}_{B}\leq\dfrac{\rho_{L}-(1+\nu)\mu}{\Phi^{-1}(\alpha)}$
This allows us to study the case in which each bank makes a unilateral decision to lend to each group, without considering the risk pooling of lending to more than one group.
Lemma 2 (Variance Threshold for Bank $L$).
@@ 3 Single-Period Setting
Consider Bank $L$, which offers loans at a normalized interest rate of $1$, lending to a group $B$ with a random payoff given by $\pi_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$. Bank $L$ lends to group $B$ in period $t$ if and only if:
$\hat{\sigma}_{BL}^{(t)}\leq\tilde{\sigma}_{L}\quad\text{Where: }\quad\tilde{\sigma}_{L}\equiv\frac{\rho_{L}-(\mu_{W}+\mu_{B}+\Phi^{-1}(\alpha)\sigma_{W})}{\Phi^{-1}(\alpha)}$
In other words, there is an upper bound $\tilde{\sigma}_{L}$ on the Bank’s beliefs about the variance of group B, such that the bank is only willing to lend to group $B$ if it believes that group has lower variance than this threshold. Otherwise, its risk of a shortfall exceeds $\alpha\%$.
A similar analysis for the subprime bank $H$ reveals that it tolerates a higher level of variance. We have:
Lemma 3 (Variance Threshold for Bank $H$).
Bank $H$ lends to group $B$ in period $t$ if and only if:
@@
$\hat{\sigma}_{BH}^{(t)}\leq\tilde{\sigma}_{H}\quad\text{Where: }\quad\tilde{\sigma}_{H}=\frac{\rho_{L}-\left[(1+\nu^{(t)})(\mu_{W}+\mu_{B})+\Phi^{-1}(\alpha)\sigma_{W}\right]}{\Phi^{-1}(\alpha)}$
@@
$0<\tilde{\sigma}_{L}<\tilde{\sigma}_{H}$
Intuitively, gains are scaled by the factor $(1+\nu^{(t)})$, providing the high-interest rate bank additional insulation against downside risk. Thus, when the perceived variance of group $i$ is between $\tilde{\sigma}_{L}$ and $\tilde{\sigma}_{H}$, the low-interest Bank $L$ will decline to lend, while the high-interest Bank $H$ can approve the loan.
4 The Subprime Trap
@@
4.1 Assumptions
We make several assumptions. These model stylized facts that characterize the subprime trap. First, we suppose that the variance of both groups is in fact below the threshold $\tilde{\sigma}_{L}$, though the variances are not necessarily identical.
Assumption 4 (Both groups are creditworthy).
$\sigma_{W}\leq\sigma_{B}<\tilde{\sigma}_{L}$
@@ 4.1 Assumptions
We characterize beliefs under imperfect information in which lending to group B does not occur. We suppose that Bank $L$ holds a prior that the variance of repayments for group $B$ is $\hat{\sigma}_{BL}^{(0)}$ such that:
@@
$\hat{\sigma}_{BL}^{(0)}>\tilde{\sigma_{L}}$
Where $\tilde{\sigma}_{L}$ is the maximum variance that Bank $L$ (the main lender) can tolerate under its constraint $\mathrm{VaR}$. Based on its initial assessment, lending violates the risk limits of Bank $L$.
Third, Bank $H$, which operates at a higher interest rate and has a correspondingly higher risk tolerance. The high-rate bank has a lending threshold $\tilde{\sigma}_{H}$ such that $\tilde{\sigma}_{H}>\tilde{\sigma}_{L}$. Denoting bank H’s prior by $\hat{\sigma}_{BH}^{(0)}$, we suppose that:
@@
$\hat{\sigma}_{BH}^{(0)}\leq\tilde{\sigma}_{H}$
@@
4.2 Equilibrium
In each period, banks choose simultaneously whether to lend to groups $W$ and $B$. Both types of bank approve group $W$’s loan application, and group $W$ chooses the lower-rate, mainstream bank. Based on their prior beliefs about B’s variances, $H$ approves $B$’s loan application, but $L$ does not. If Bank $L$ does not lend to group $B$, then group $B$ is left with the subprime option from Bank $H$. But since Bank $L$’s belief update depends on observing a return from group $B$, if it does not lend to group $B$ at period $t$, it does not have an updated posterior belief to use as its prior at period $t+1$: its beliefs do not change. The crucial observation is that if Bank $L$ persistently withholds loans to group $B$, it never observes the repayment data needed to update its inflated variance estimate, and its belief remains at $\hat{\sigma}_{BL}^{(0)}$.
@@ 4.2 Equilibrium
Suppose Assumptions 1-6 hold. Then, there exists a Bayesian subgame-perfect equilibrium in which, in every period, Bank $L$ lends exclusively to group $W$, and Bank $H$ lends to group $B$ with high probability, permanently relegating group $B$ to subprime loans.
@@
First, in the base case, the bank does not lend to group B. We have that $A_{ij}^{(1)}=0$, since, by Lemma 1, if the bank’s prior belief is $\hat{\sigma}_{BL}^{(0)}$ and $\hat{\sigma}_{BL}^{(0)}>\tilde{\sigma}_{L}$, then $A_{ij}^{(1)}=0$.
To show the induction step, we show that if the bank does not lend to B in period $t$, then it does not lend to B in period $t+1$. For this, we have that $A_{ij}^{(t)}=0\implies\hat{\sigma}_{ij}^{(t+1)}=\hat{\sigma}_{ij}^{(t)}$, because no updating occurs, but since $A_{ij}^{(t+1)}=1$ if and only if $\hat{\sigma}_{ij}^{(t+1)}\leq\tilde{\sigma}_{L}$, we have that $\hat{\sigma}_{B}^{(t)}>\tilde{\sigma}_{L}$ and $\hat{\sigma}_{ij}^{(t+1)}=\hat{\sigma}_{ij}^{(t)}\implies\hat{\sigma}_{ij}^{(t+1)}>\tilde{L}$, so that $A_{ij}^{(t+1)}=0$. This verifies the induction step.
@@
Next, consider Bank $H$. By hypothesis, Bank $H$ believes that $\sigma_{BH}^{(0)}\leq\tilde{\sigma}_{H}$, and therefore lending to group $B$ satisfies its risk constraint in the initial period. Moreover, since the true mean $\mu_{B}$ is positive, lending to group $B$ is expected to yield a positive return. Bank $H$ lends to group $B$ with high probability, because in each stage there is a (vanishing) probability of observing a payoff realization that increases Bank $H$’s posterior belief about group B’s variance above its threshold. This probability goes to zero by the SLLN, however.
@@
Finally, note that because Bank $L$ never lends to group $B$, it never observes any repayment outcomes from that group. This implies that the variance estimate for group $B$ remains at $\sigma_{BL}^{(0)}>\tilde{\sigma}_{L}$ in every period. As a consequence, there is no incentive for Bank $L$ to deviate from its strategy of lending only to group $W$, and similarly, neither Bank $H$ nor the borrower groups have incentives to deviate from their prescribed actions. Hence, the described strategy profile constitutes a subgame-perfect equilibrium. ∎
@@
Even if the true variance of group $B$ were such that $\sigma_{B}<\tilde{\sigma}_{L}$, Bank $L$ would remain unaware of this fact unless it extends credit to group $B$. The absence of new data prevents Bank $L$ from updating its risk assessment, leading to an inefficient outcome in which group $B$ continues to access only subprime credit.
@@
5 Escaping the Subprime Trap via Subsidies
We now show that a subsidy or partial lending guarantee can help group $B$ escape the subprime trap. The key idea is to cover sufficient downside risk so that Bank $L$ is induced to lend to group $B$. When this occurs, Bank $L$ can gather information on group $B$’s repayment performance. Since, by Assumption 1, group $B$, is in fact creditworthy, Bank $L$’s beliefs will eventually converge to group $B$’s actual variance, which, by Assumption 2, is low enough that $L$’s VaR constraint will be satisfied. Once enough observations have been accumulated, Bank $L$ will update its risk assessment so that the estimated variance satisfies $\hat{\sigma}_{BL}^{(t)}<\tilde{\sigma}_{L}$, eliminating the need for further subsidy.
@@ 5 Escaping the Subprime Trap via Subsidies
${{\hat{\sigma}_{i}}}^{(m)}\searrow\sigma_{B}$ almost surely as $m\to\infty$
By Proposition 4.1, this is the condition required for Bank L to lend to group $B$, so that the $\mathrm{VaR}$ constraint is satisfied. In other words, if the mainstream bank were to lend to group $B$ for a sufficiently long period of time, it would learn about the variance of returns due to lending to group $B$. The problem is then how to induce the bank to do so.
@@
${s^{*}}^{(t)}=\inf\left\{s\geq 0\>:\>\mathbb{P}\left(\Pi_{j}^{(t)}+s<\rho_{j}\right)\leq\alpha\right\}$
@@
${s^{*}}^{(t)}=\max\Bigl{\{}0,\;\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BL}^{(t)}+\sigma_{W})\Bigr{\}}$
@@
Parameters: $\rho_{L}$, $\alpha$, $\mu$, initial variance $\hat{\sigma}_{BL}^{(0)}$, horizon $T$.
Set $m\leftarrow 0$  (# of observations from group $B$)
for $t\in 1,\cdots,T$ do
     while ${\hat{\sigma}_{BL}}^{(t)}>{\tilde{\sigma}_{BL}}$ do
         if ${A_{BL}}^{(t-1)}=1$ then
              $\>\quad{\hat{\sigma}_{BL}}^{(t)}\leftarrow\left[\dfrac{1}{\sum_{m<t}\mathbb{I}\left\{{A_{BL}}^{(m)}=1\right\}+1}\left({\hat{\sigma}_{B}}^{2(0)}+\sum_{m<t}\mathbb{I}\left\{{A_{BL}}^{(m)}=1\right\}\Bigl{(}{\pi_{BL}}^{(m)}-{\overline{\pi}_{BL}}^{(t-1)}\Bigr{)}^{2}\right)\right]^{1/2}$
@@
         end if
         ${s^{*}}^{(t)}\leftarrow\max\Bigl{\{}0,\;\rho_{L}-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BL}^{(t)}+\sigma_{W})\Bigr{\}}$
     end while
end for
The subsidy incentivizes Bank $L$ to lend to group $B$, generating observations that allow its risk assessment to converge to the true variance $\sigma_{B}^{2}$. Once the updated estimate satisfies $\hat{\sigma}_{BL}^{(t)}<\tilde{\sigma}_{L}$, it follows from Lemma 2 that Bank $L$ meets its $\mathrm{VaR}$ constraint without subsidy (i.e., ${s^{*}}^{(t)}$), and no further external support is required.
@@
Assume that the true distribution of repayments for group $B$ is $\mathcal{N}(\mu,\sigma_{B}^{2})$ with $\sigma_{B}<\tilde{\sigma}_{L}$, but that Bank $L$ initially believes that the variance is $\hat{\sigma}_{BL}^{(0)}>\tilde{\sigma}_{L}$. Under the subsidy mechanism described above (Algorithm 1), with probability one there exists a finite time $\tau$ such that for all $t\geq\tau$, Bank $L$ updates its variance estimate so that $\hat{\sigma}_{BL}^{(t)}\to\sigma_{B}<\tilde{\sigma}_{L}$ and consequently ${s^{*}}^{(t)}$. Beyond time $\tau$, Bank $L$ continues lending to group $B$ without further subsidy.
In short, the subsidy induces the bank to update their information about minority group creditworthiness. It is also temporary when there exist otherwise creditworthy borrowers trapped in subprime loans: subsidizing information discovery helps banks to learn the true variance of lender groups.
@@
For $t>\tau,\>\nu^{(t)}=0$.
In any round in which the low-interest bank would approve group $B$’s loan, the high-interest bank will choose to set its premium to zero. If it chooses any positive premium, neither group will accept the loan. If it chooses a premium of zero, both groups will randomize over accepting the loan from bank $B$ and accepting the loan from bank $H$. Since this persists for all $t>\tau$, this reduces interest rates permanently – competition reduces the premiums charged by subprime lenders.
6 Conclusion
Inequalities in algorithmic lending practices can persist even when banks believe minority applicants to be as creditworthy as white applicants. We study one explanation for this phenomenon: the role of risk-management constraints, specifically $\mathrm{VaR}$, in contributing to persistent disparities in lending. We have shown that risk management constraints can lead lenders to refuse loans even with positive Net Present Value. This forces minority applicants to accept high-interest rate loans, with negative implications for their financial prospects, a situation we describe as the ”subprime trap” equilibrium. We emphasize the role of inaccurate prior beliefs about the risk of lending to minority groups, which leads to systematic exclusion and higher borrowing costs.
We showed that targeted, temporary interventions, such as subsidies or guarantees, can break this cycle by allowing banks to learn true risk profiles. These findings suggest practical avenues for addressing lending disparities and offer a framework to explain failures of fairness in algorithmic lending decisions.
Appendix

### v2

@@ Introduction
If bank $j$ lends to group $i$, it observes a return $\pi_{it}$ and then updates its posterior beliefs about the variance $\sigma_{i}^{2}$. Suppose that the bank observes a sequence of returns $\{\pi_{1},\pi_{2},\dots,\pi_{M}\}$. Then, it has posterior belief about the variance of group $i$:
$\sigma_{B}^{2}|\pi_{1},\dots,\pi_{M}\sim$
$\text{Inv-}\Gamma\left(a_{0}+\dfrac{np_{B}+M}{2},b_{0}+\dfrac{np_{B}+\sum_{m=1}^{M}(\pi_{Bm}-\mu_{B})^{2}}{2}\right)$
So that:
$\mathbb{E}[\sigma_{B}^{2}|\pi_{1},\dots,\pi_{M}]=\dfrac{b_{0}+\dfrac{np_{B}+\sum_{m=1}^{M}(\pi_{Bm}-\mu_{B})^{2}}{2}}{a_{0}+\dfrac{np_{B}+M}{2}-1}$
In other words, the bank’s posterior belief about the variance of the minority group depends both on the depth of historical credit files ($np_{B}$) and the observed variance of repayments ($\sum_{m=1}^{M}(\pi_{Bm}-\mu_{B})^{2}$), conditional on lending to the minority group.
@@
Our Modeling Choices
VaR and CVaR/ES Value-at-Risk is a commonly used risk management measure in practice, and is explicitly referenced in the Basel III banking regulations (Li 2014; Chang et al. 2019). We could also study the Conditional Value-at-Risk, or Expected Shortfall, which is related to VaR via the formula: $ES(X)=\frac{1}{\alpha}\int_{0}^{\alpha}VaR_{\phi}(X)\>d\phi$. The intuitions underlying our model carry through to the CVaR/ES risk metric: there is a threshold variance at which the CVaR/ES will be violated. We show that the CVaR/ES threshold is more conservative than the VaR risk metric, and hence requires a larger subsidy to escape the subprime trap, in the Appendix.
@@
Inaccurate prior beliefs Our model supposes that banks have inaccurate beliefs about the creditworthiness of applicants. This is a plausible outcome when decisions are made on the basis of credit scoring data that is either biased or noisy. Credit scores may be inherently noisier representations of underlying default risk for minority groups (Blattner and Nelson 2021a), while the components of the score may themselves be low quality indicators of repayment ability (Rice and Swesnik 2013). Empirical work shows that banks are more likely to avoid lending to minority neighborhoods, thereby lowering the quality of information received about lenders in those neighborhoods (Blattner and Nelson 2021b). Noisier estimates of credit risk are likely to lead to inflated prior estimates of variance in practice.
Risk-pooling We study three cases: where banks lend unilaterally to each group, and where banks lend to both groups. In the unilateral setting, there is no risk pooling, and banks make decisions based only on single group characteristics. However, when banks are willing to lend to both groups, risk pooling occurs, where the lower-risk group essentially subsidizes the risk of lending to the higher-risk group. This case is studied in (Donahue and Barocas 2021). Below, we derive variance thresholds for each of the three cases.
Single-Period Setting
We first study the stage game, to show how the VaR constraint affects lending to group $B$ in each period. Note first that we assume that Bank H charges an interest rate premium to derive variance thresholds. We then prove that this occurs. Second, note that we derive variance thresholds for the pooled variance – thresholds under which banks will lend to both groups. Because the variance thresholds are ordered, it follows that if the bank’s variance pooled threshold is not satisfied, it will not unilaterally lend to the minority group. It also follows that if the bank’s pooled threshold is satisfied, the bank will unilaterally lend to the majority group.
First we assume that Banks have unilateral thresholds for lending to each group:
Lemma 2 (Unilateral Lending Thresholds).
$\tilde{\sigma}^{2}_{L}$ $=\left(\dfrac{\rho_{L}-\mu}{\Phi^{-1}(\alpha)}\right)^{2}$
$\tilde{\sigma}^{2}_{H}$ $=\left(\dfrac{\rho_{L}-(1+\nu)\mu}{\Phi^{-1}(\alpha)}\right)^{2}$
These represent the thresholds under which each bank is willing to lend unilaterally to a given group.
Each bank also has a pooled variance threshold, under which it is willing to lend to both groups.
Lemma 3 (Pooled Variance Threshold for Bank $L$).
@@
Consider Bank $L$, which offers loans at a normalized interest rate of $1$, lending to group $B$ with a random payoff given by $\pi_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$. Bank $L$ lends to both groups in period $t$ if and only if group B’s variance satisfies:
@@
$\hat{\sigma}^{2}_{BLt}\leq\tilde{\sigma}_{L}^{2,\textrm{pool}}\equiv\left(\frac{\rho_{L}-(\mu_{W}+\mu_{B}+\Phi^{-1}(\alpha)\sigma_{W})}{\Phi^{-1}(\alpha)}\right)^{2}$
In other words, there is an upper bound $\tilde{\sigma}^{L}$ on the Bank’s beliefs about the variance of group B, such that the bank is only willing to lend to group $B$ if it believes that group has lower variance than this threshold. Otherwise, its risk of a shortfall exceeds $\alpha\%$.
A similar analysis for the subprime bank $H$ shows that it tolerates a higher level of variance. We have:
Lemma 4 (Pooled Variance Threshold for Bank $H$).
Bank $H$ lends to group $B$ and group $W$ in period $t$ if and only if $\hat{\sigma}^{2}_{BHt}\leq\tilde{\sigma}^{H}_{\text{pool}}$, where:
$\tilde{\sigma}_{H}^{2,\text{pool}}=\left(\frac{\rho_{H}-\left[(1+\nu_{t})(\mu_{W}+\mu_{B})+\Phi^{-1}(\alpha)\sigma_{W}\right]}{\Phi^{-1}(\alpha)}\right)^{2}$
@@
While $\nu_{t}>0$, and for $\alpha<.1$:
$0<\tilde{\sigma}^{2}_{L}<\tilde{\sigma}^{2,\text{pool}}_{L}<\tilde{\sigma}^{2}_{H}<\tilde{\sigma}_{H}^{2,\text{pool}}$
Intuitively, gains are scaled by the factor $(1+\nu_{t})$, providing the high-interest rate bank insulation against downside risk. Thus, when the perceived variance of group $i$ is between $\tilde{\sigma}^{2}_{L}$ and $\tilde{\sigma}^{2}_{H}$, the low-interest Bank $L$ will decline to lend, while the high-interest Bank $H$ can approve the loan.
The Subprime Trap
@@
Assumptions
We make several assumptions. These model stylized facts that characterize the subprime trap. First, we suppose that the variance of both groups is in fact below the threshold $\tilde{\sigma}^{2}_{L}$, though the variances are not necessarily identical.
Assumption 3 (Both groups are in fact creditworthy).
$\sigma^{2}_{W}\leq\sigma^{2}_{B}<\tilde{\sigma}^{2}_{L}$
@@
We characterize beliefs under imperfect information in which lending to group B does not occur. We suppose that Bank $L$ holds a prior that the variance of repayments for group $B$ is $\hat{\sigma}^{2}_{BL,0}$ such that:
@@
$\hat{\sigma}^{2}_{BL,0}>\tilde{\sigma}^{2,\text{pool}}_{L}$
Where $\tilde{\sigma}^{2,\text{pool}}_{L}$ is the maximum variance that Bank $L$ (the main lender) can tolerate under its constraint $\mathrm{VaR}$. Based on its initial assessment, lending violates the risk limits of Bank $L$.
Third, Bank $H$, which operates at a higher interest rate and has a correspondingly higher risk tolerance. The high-rate bank has a lending threshold $\tilde{\sigma}^{2}_{H}$ such that $\tilde{\sigma}^{2}_{H}>\tilde{\sigma}^{2,\text{pool}}_{L}$. Denoting bank H’s prior by $\hat{\sigma}^{2}_{BH,0}$, we suppose that:
@@
$\hat{\sigma}^{2}_{BH,0}\leq\tilde{\sigma}^{2}_{H}$
@@
Equilibrium
In each period, banks choose simultaneously whether to lend to groups $W$ and $B$. Both types of bank approve group $W$’s loan application, and group $W$ chooses the lower-rate, mainstream bank. Based on their prior beliefs about B’s variances, $H$ approves $B$’s loan application, but $L$ does not. If Bank $L$ does not lend to group $B$, then group $B$ is left with the subprime option from Bank $H$. But since Bank $L$’s belief update depends on observing a return from group $B$, if it does not lend to group $B$ at period $t$, it does not have an updated posterior belief to use as its prior at period $t+1$: its beliefs do not change. The crucial observation is that if Bank $L$ persistently withholds loans to group $B$, it never observes the repayment data needed to update its inflated variance estimate, and its belief remains at $\hat{\sigma}_{BL,0}^{2}$.
@@
Suppose Assumptions 1-5 hold. Then, there exists a Bayesian subgame-perfect equilibrium in which, in every period, Bank $L$ lends exclusively to group $W$, and Bank $H$ lends to group $B$ with high probability, permanently relegating group $B$ to subprime loans.
@@
First, in the base case, the bank does not lend to group B. We have that $A_{ij1}=0$, since, by Lemma 1, if the bank’s prior belief is $\hat{\sigma}^{2}_{BL0}$ and $\hat{\sigma}^{2}_{BL0}>\tilde{\sigma}^{2,\text{pool}}_{L}$, so that the VaR constraint is violated, by Lemma 1, and hence lending does not occur in period 1.
To show the induction step, we show that if the bank does not lend to B in period $t$, then it does not lend to B in period $t+1$. For this, we have that $A_{ijt}=0\implies\hat{\sigma}^{2}_{ijt+1}=\hat{\sigma}^{2}_{ijt}$, because no updating occurs. but since $A_{ijt+1}=1$ if and only if $\hat{\sigma}_{ijt+1}\leq\tilde{\sigma}^{2,\text{pool}}_{L}$, we have that $\hat{\sigma}^{2}_{Bt}>\tilde{\sigma}^{2,\text{pool}}_{L}$ and $\hat{\sigma}^{2}_{ijt+1}=\hat{\sigma}_{ijt}^{2}\implies\hat{\sigma}^{2}_{ijt+1}>\tilde{\sigma}^{2}_{L}$, so that $A_{ijt+1}=0$. This verifies the induction step.
@@
Next, consider Bank $H$. By hypothesis, Bank $H$ believes that $\sigma^{2}_{BH0}\leq\tilde{\sigma}^{2}_{H}$, and therefore lending to group $B$ satisfies its risk constraint in the initial period. Moreover, since the true mean $\mu_{B}$ is positive, lending to group $B$ is expected to yield a positive return. Bank $H$ lends to group $B$ with high probability, because in each stage there is a (vanishing) probability of observing a payoff realization that increases Bank $H$’s posterior belief about group B’s variance above its threshold. This probability goes to zero by the SLLN, however.
@@
Finally, note that because Bank $L$ never lends to group $B$, it never observes any repayment outcomes from that group. This implies that the variance estimate for group $B$ remains at $\hat{\sigma}^{2}_{BL0}>\tilde{\sigma}^{2,\text{pool}}_{L}$ in every period. As a consequence, there is no incentive for Bank $L$ to deviate from its strategy of lending only to group $W$, and similarly, neither Bank $H$ nor the borrower groups have incentives to deviate from their prescribed actions. Hence, the described strategy profile constitutes a subgame-perfect equilibrium. ∎
@@
Even though the true variance of group $B$ is such that $\sigma^{2}_{B}<\tilde{\sigma}^{2,\text{pool}}_{L}$, Bank $L$ remains unaware of this fact unless it extends credit to group $B$. The absence of new data prevents Bank $L$ from updating its risk assessment, leading to an inefficient outcome in which group $B$ continues to access only subprime credit.
@@
Escaping the Subprime Trap via Subsidies
We now show that a subsidy or partial lending guarantee can help group $B$ escape the subprime trap. The key idea is to cover sufficient downside risk so that Bank $L$ is induced to lend to group $B$. When this occurs, Bank $L$ can gather information on group $B$’s repayment performance. Since, by Assumption 1, group $B$, is in fact creditworthy, Bank $L$’s beliefs will eventually converge to group $B$’s actual variance, which, by Assumption 2, is low enough that $L$’s VaR constraint will be satisfied. Once enough observations have been accumulated, Bank $L$ will update its risk assessment so that the estimated variance satisfies $\hat{\sigma}^{2}_{BLt}<\tilde{\sigma}^{2}_{L}$, eliminating the need for further subsidy.
@@
${{\hat{\sigma}^{2}_{BLm}}}\to\sigma^{2}_{B}$ almost surely as $m\to\infty$
By Lemma 3, this is the condition required for Bank L to lend to group $B$, so that the $\mathrm{VaR}$ constraint is satisfied. In other words, if the mainstream bank were to lend to group $B$ for a sufficiently long period of time, it would learn about the variance of returns due to lending to group $B$. The problem is then how to induce the bank to do so.
@@
${s_{t}^{*}}=\inf\left\{s\geq 0\>:\>\mathbb{P}\left(\Pi_{jt}+s<\rho\right)\leq\alpha\right\}$
@@
${s_{t}^{*}}=\max\Bigl\{0,\;\rho-(\mu_{W}+\mu_{B})-\Phi^{-1}(\alpha)\,(\hat{\sigma}_{BLt}+\sigma_{W})\Bigr\}$
@@
1:Parameters: $\rho$, $\alpha$, $\mu$, prior parameters $(a_{0},b_{0})$, $np_{B}$, threshold $\tilde{\sigma}^{2}_{L}$, horizon $T$
2:Initialize:
@@
[…]
6:for $t\in 1,\cdots,T$ do
7:  $\hat{\sigma}^{2}_{BLt}\leftarrow\frac{b_{t}}{a_{t}-1}$ $\triangleright$ Current estimate
8:  if $\hat{\sigma}^{2}_{BLt}>\tilde{\sigma}^{*}_{L}$ then
9:   $s_{t}\leftarrow s^{*}_{t}(\hat{\sigma}^{2}_{BLt})$
10:   Bank L lends to B with subsidy $s_{t}$
11:   Observe return $\pi_{B,t}$
12:   $M\leftarrow M+1$
13:   $a_{t}\leftarrow a_{0}+\frac{np_{B}+M}{2}$
14:   $b_{t}\leftarrow b_{0}+\frac{np_{B}+\sum_{m=1}^{M}(\pi_{B,m}-\mu)^{2}}{2}$
15:  else
16:   No subsidy needed.
17:  end if
18:end for
The subsidy incentivizes Bank $L$ to lend to group $B$, generating observations that allow its risk assessment to converge to the true variance $\sigma_{B}^{2}$. Once the updated estimate satisfies $\hat{\sigma}^{2}_{BLt}<\tilde{\sigma}^{2}_{L}$, it follows from Lemma 2 that Bank $L$ meets its $\mathrm{VaR}$ constraint without subsidy (i.e., $s_{t}^{*}$}), and no further external support is required.
@@
Assume that the true distribution of repayments for group $B$ is $\mathcal{N}(\mu,\sigma_{B}^{2})$ with $\sigma^{2}_{B}<\tilde{\sigma}^{2}_{L}$, but that Bank $L$ initially believes that the variance is $\hat{\sigma}^{2}_{BL0}>\tilde{\sigma}^{2}_{L}$. Under the subsidy mechanism described above (Algorithm 1), with probability one there exists a finite time $\tau$ such that for all $t\geq\tau$, Bank $L$ updates its variance estimate so that $\hat{\sigma}^{2}_{BLt>\tau}\to\sigma^{2}_{B}<\tilde{\sigma}^{L}$ and consequently ${s_{t}^{*}}\to 0$. Beyond time $\tau$, Bank $L$ continues lending to group $B$ without further subsidy.
In short, the subsidy induces the bank to update their information about minority group creditworthiness. It is also temporary when there exist otherwise creditworthy borrowers trapped in subprime loans: subsidizing information discovery helps banks to learn the true variance of lender groups. Further, while algorithmic audits could correct bias, subsidies can avoid privacy hurdles and scale faster (Cowgill and Tucker 2019).
@@
For $t>\tau,\>\nu_{t}=0$.
In any round in which the low-interest bank would approve group $B$’s loan, the high-interest bank will choose to set its premium to zero. If it chooses any positive premium, neither group will accept the loan. If it chooses a premium of zero, both groups will randomize over accepting the loan from bank $B$ and accepting the loan from bank $H$. Since this persists for all $t>\tau$, this reduces interest rates permanently: competition reduces the premiums charged by subprime lenders.
Conclusion
Inequalities in algorithmic lending practices can persist even when minority applicants are as creditworthy as white applicants. We study one explanation for this phenomenon: the role of risk-management constraints, specifically $\mathrm{VaR}$, in contributing to persistent disparities in lending. We have shown that risk management constraints can lead lenders to refuse loans even with positive Net Present Value. This forces minority applicants to accept high-interest rate loans, with negative implications for their financial prospects, a situation we describe as the “subprime trap” equilibrium. We emphasize the role of inaccurate prior beliefs about the risk of lending to minority groups, which leads to systematic exclusion and higher borrowing costs.
Our theoretical results align with documented patterns in mortgage lending markets. First, the persistence of our subprime trap equilibrium mirrors the decade-long stability of racial lending disparities documented in mortgage markets, even as algorithmic lending has expanded (Bartlett et al. 2022). Second, our model’s prediction that mainstream banks never update their beliefs about minority creditworthiness is consistent with (Bayer, Ferreira, and Ross 2016). Third, the equilibrium’s existence despite equal creditworthiness directly explains the puzzle identified by (Popick 2022) and (Crosignani and Le 2023): why minorities with comparable credit scores remain concentrated at high-cost lenders.
Implementation of our subsidy mechanism maps naturally onto existing policy infrastructure. The Federal Housing Administration’s mortgage insurance programs already function as partial guarantees, covering lender losses when borrowers default on FHA-insured loans (U.S. Department of Housing and Urban Development 2024). Similarly, the Community Reinvestment Act provides regulatory credits to banks for community development loans and investments in low- and moderate-income neighborhoods (U.S. Congress 2013; Board of Governors of the Federal Reserve System 2024). Our analysis suggests these programs could achieve greater impact if restructured to explicitly target variance reduction during a learning period; guaranteeing tail risks for a finite horizon would enable banks to update their risk assessments while avoiding solvency-affecting losses.
Our analysis focuses on one specific mechanism: variance-based discrimination through VaR constraints. Observed lending disparities likely result from multiple interacting factors including historical discrimination, geographic segregation, and differential access to financial services.
Nonetheless, targeted, temporary interventions, such as subsidies or guarantees, can break this cycle by allowing banks to learn true risk profiles. These findings suggest practical avenues for addressing lending disparities and offer a framework to explain failures of fairness in algorithmic lending decisions.
Appendix A Appendix